# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ 3.2. Lagrangian chaos for stochastic PEs
  • Let $\mu$ be the unique invariant measure for (1.1) on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$ given by Theorem 3.3. The unique invariant measure on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$ for the Lagrangian process $(v_{t},\mathbf{x}_{t})$ is given by $\mu\otimes\mathrm{Leb}$.
  • The projective processes $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ and $(v_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})_{t}$ have a unique invariant measures on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}\times\mathrm{Pr}_{2}$.
Proof of Corollary 3.13.
Let us first consider the Lagrangian process. The invariance of $\mu\otimes\mathrm{Leb}$ is immediate due to the incompressibility of $u_{t}$. By (3.2) and the Krein-Milman theorem, it is enough to show the uniqueness of ergodic measures. The latter follows from a well-known generalization of Khasminskii’s uniqueness result (see, for example, [25, Proposition 4.1.1]) and Propositions 3.11-3.12.
The case of projective processes is similar. It is enough to recall that from the first part of the proof, we have for all invariant measure $\nu$ of the projective process we have $\nu(\cdot\times\mathrm{Pr}_{2})=\mu\otimes\mathrm{Leb}$, and therefore existence follows from Krylov-Bogoliubov (see [16, Lemma 3.15]). The rest of the argument proceeds as in the Lagrangian process case. ∎
@@
We begin by proving (3.4) (resp., (3.5)). The latter follows from Propositions 3.8-3.10 and Theorem 3.3 by applying combining the multiplicative ergodic theorem [11, Theorem 3.4.1] (resp., random multiplicative ergodic theorem [50, Theorem III.1.2]) and [16, Proposition 4.17] to the linear cocycle $\mathcal{A}_{\omega,(v,\mathbf{x})}^{t}$ induced by (1.1). Let us note that multiplicative ergodic theorems ensure the existence of the limits (3.4) and (3.5) only for $\mu\otimes\mathrm{Leb}$-a.a. $(v,x)\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$ (here we also used Theorem 3.3). As noticed in [16, Remark 2.2], the strong Feller property as in Proposition 3.11 allow to upgrade the statement to for all $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$. For the reader’s convenience, we include some details. We only comment on the case (3.4) as the other is similar. Suppose that (3.4) holds for a.a. $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$. Let =$\mathcal{L}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{(\omega,v,\mathbf{x})\in\Omega\times\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}\,:\,\lim_{t\to\infty}t^{-1}\log|\nabla\phi^{t}_{\omega,v}(\mathbf{x})|=\lambda\}$ and set
$\psi_{\mathcal{L}}(v,\mathbf{x})\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathbf{E}[\mathbf{1}_{\mathcal{L}}(\cdot,v,\mathbf{x})]=\mathbf{E}[\mathbf{1}_{\mathcal{L}}(\theta^{t}\cdot,v,\mathbf{x})],$
@@
for all $t>0$, where the last equality follows from $\mathbf{P}\cdot(\theta^{t})^{-1}=\mathbf{P}$. From the assumption it follows that $\psi_{\mathcal{L}}(v,\mathbf{x})=1$ for $\mu\otimes\mathrm{Leb}$-a.a. $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$. By the latter and invariance of $\mu\otimes\mathrm{Leb}$ for the Lagrangian process (with $L_{t}$ as in (3.22)),
${1=\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}}\psi_{\mathcal{L}}(v,\mathbf{x})\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}=\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}}L_{t}\psi_{\mathcal{L}}(v,\mathbf{x})\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}}$
@@
for all $t>0$ (in the following, we pick $t=1$). Hence, $L_{1}\psi_{\mathcal{L}}(v,\mathbf{x})=1$ for $\mu\otimes\mathrm{Leb}$-a.a. $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$, and by strong Feller property of Proposition 3.11, we have $L_{1}\psi_{\mathcal{L}}(v,\mathbf{x})=1$ for all $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}$. Now, the independence of $\theta^{1}\omega$ and $(u_{1}(\omega),\mathbf{x}_{1}(\omega))$ yield
${\int_{\Omega}\mathbf{1}_{\mathcal{L}}(\theta^{1}\omega,v_{1}(\omega),\mathbf{x}_{1}(\omega))\,\mathrm{d}\mathbf{P}(\omega)=L_{1}\psi_{\mathcal{L}}(v,\mathbf{x})=1}.$
@@
Thus, $\mathbf{1}_{\mathcal{L}}(\theta^{1}\cdot,v_{1},\mathbf{x}_{1})$ a.s. Hence, the conclusion follows by noticing that $(\omega,v,\mathbf{x})\in\mathcal{L}$ if and only if $(\theta^{1}\omega,v_{1}(\omega),\mathbf{x}_{1}(\omega))\in\mathcal{L}$ due to the invertibility of $\nabla\phi^{t}_{\cdot,\cdot}(\cdot)$.
@@ 4. Integrability of invariant measures
In this section, we prove Proposition 3.9. As discussed in Subsection 3.3, Proposition 3.9 was used to obtain the existence of the (top) Lyapunov exponent in Theorems 1.1 and 3.5. Actually, we prove a more refined result compared to Proposition 3.9. Indeed, we obtain an $L^{p}$-integrability of $\nabla u$ with respect to $\mu\otimes\mathrm{Leb}$ where $\mu$ is an invariant measure for the stochastic PEs (1.1). Moreover, we can prove the latter under minimal requirements on $Q$. More precisely, in this subsection, we work under the following assumption.
@@
The main result of this section reads as follows.
@@
(4.1) $\int_{\mathbb{H}_{0}^{1}(\mathbb{T}^{3})}\int_{\mathbb{T}^{3}}|\nabla u(\mathbf{x})|^{p}\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)\leq C(p,\mathcal{E})$
@@
(4.2) $\int_{\mathbb{H}_{0}^{1}(\mathbb{T}^{3})}\int_{\mathbb{T}^{3}}|\nabla^{2}v(\mathbf{x})|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)$
@@
$\lesssim\|\nabla v\|_{L^{2}(\mathbb{T}^{3})}^{2},$
where the last term can be estimated by means of the energy inequality. From the estimate (4.7), one sees that we have at our disposal more integrability in the vertical direction than in the horizontal ones. The above inequality, along with Proposition 4.3, suggests the validity of (4.5) with $p$ close to $1$.
@@ 4.1. Auxiliary results
In this subsection, we collect some useful results in the proof of Theorem 4.2 and Proposition 4.4. We begin with an elementary result concerning the integral operator on the one-dimensional torus $\mathbb{T}=[0,1]$. For $f\in L^{1}(\mathbb{T})$, let
@@
In the proof below, we use the notion of the $H^{\infty}$-calculus and $\mathcal{R}$-sectoriality for various versions of the Laplace operator. We refer to either [48, Chapter 10] or [66, Chapters 3 and 4] for details and notation.
We begin by proving (4.10). Let $A\stackrel{{\scriptstyle{\rm def}}}{{=}}-\Delta_{x,y}$ and $B\stackrel{{\scriptstyle{\rm def}}}{{=}}-\partial^{2}_{z}$ be the realization of the (minus) Laplace operator on $H^{s_{0},p_{0}}_{x,y}(H^{s_{1},p_{1}}_{z})$ with domain
@@
Moreover, $A$ and $B$ have $H^{\infty}$-calculus of angle 0 by the periodic version of [48, Theorem 10.2.25]. In particular the operator $A+B$ on $\mathsf{D}(A)\cap\mathsf{D}(B)$ is $\mathcal{R}$-sectorial of angle 00 by [66, Corollary 4.5.9]. The maximal $L^{q}$-regularity estimate (4.10) now follows from the Weis’ characterization, see e.g. [66, Theorem 4.4.4] or [71].
Due to (4.10), the semigroup $(e^{t\Delta})_{t\geq 0}$ is analytic on $\mathcal{X}$ by [66, Proposition 3.5.2(i)]. Hence (4.11) follows from Young’s convolution inequality.
It remains to prove the last assertion of Lemma 4.6. The embedding essentially uses mixed-derivative estimates. Indeed, by [66, Corollary 4.5.11] for all $\theta\in(0,1)$,
@@
Before going into the proof of Proposition 4.4, let us state the following result that is a straightforward consequence of the definitions in (2.5) and will be used below without further mentioning it.
@@
The above is known to experts. However, for the reader’s convenience, we include a short proof based on a duality argument. Note that the condition $-d>r-\frac{d}{q}$ is equivalent to $-r-\frac{d}{q^{\prime}}>0$ where $q^{\prime}=\frac{q}{q-1}$. By the vector-valued Sobolev embeddings (see e.g. [60, Proposition 7.4]) we have
@@ 5.1.1. Sobolev estimates for the PEs nonlinearity
Let $\eta,\xi$ be stopping times such that $0\leq\eta\leq\xi\leq\tau_{j}$ a.s. for some $j\geq 1$. By well-known $L^{2}$-estimates for the stochastic heat equation, there exists a constant $C_{0}>0$ independent of $(j,\eta,\xi)$ such that (see e.g., [10, Lemma4.1])
@@
Recall that $\mathcal{E}_{\sigma}=\|Q\|_{\mathscr{L}_{2}(\mathbb{L}^{2}_{0},\mathbb{H}^{\sigma}_{0})}$ and fix $\sigma^{\prime}<1+\sigma$. We claim that, if there exists $\eta\in[2,\sigma^{\prime}]$ and $\gamma\in{\mathbb{N}}_{\geq 1}$ for which
@@
then there exists $\gamma^{\prime}$ depending only on $\gamma,\sigma$ and $\sigma^{\prime}$ such that
@@
(5.10) $\int_{\mathbb{H}^{1}_{0}(\mathbb{T}^{3})}\log^{[\gamma^{\prime}]}(R_{\gamma}^{\prime}(1+\|v\|_{H^{\eta^{\prime}}}))\,\mathrm{d}\mu(v)\lesssim_{\mathcal{E}_{\sigma}}1\ \text{ where }\ \eta^{\prime}\stackrel{{\scriptstyle{\rm def}}}{{=}}(\eta+\tfrac{1}{2})\wedge\sigma^{\prime}$
and $R^{\prime}_{\gamma}$ depends only on $\sigma^{\prime},\sigma$ and $R_{\gamma}$.
@@
To prove (5.9$\Rightarrow$(5.10), we employ weighted maximal $L^{p}$-regularity estimates, see e.g., [66, Section 3.5.2] or [46, Subsection 17.2.e]. Here, we employ the weighted maximal $L^{p}$-regularity estimates for the Laplace operator on $-\Delta:H^{\eta^{\prime}+1}\subseteq H^{\eta^{\prime}-1}\to H^{\eta^{\prime}-1}$ which holds by to standard Fourier methods (for instance, from the periodic version of [48, Example 10.1.5]). Let $p\in(2,\infty)$ and $\kappa\in[0,\frac{p}{2}-1)$ be such that $\eta+1-2\frac{1}{p}>\eta^{\prime}$ and $\eta^{\prime}+1-2\frac{1+\kappa}{p}<\eta$. Note that the latter choice is always possible by using $\eta^{\prime}<\eta+\frac{1}{2}$ and choosing $\kappa$ close to $p-1$. Hence, by either [46, Corollary 17.2.37] or [66, Theorem 3.4.8], the maximal $L^{p}$-regularity of $-\Delta$ on $H^{\eta^{\prime}-1}$ with time weight $t^{\kappa}$ applied to $v^{\prime}_{t}=v_{t}-\Gamma_{t}$ (here $\Gamma_{t}=\int_{0}^{t}e^{\Delta(t-s)}Q\,\mathrm{d}W_{s}$), we obtain, a.s.,
@@
$\|v_{1}^{\prime}\|_{H^{\eta^{\prime}}}^{p}$ $\textstyle\lesssim\|v\|_{H^{\eta}}^{p}+\int_{0}^{1}t^{\kappa}\|b(v_{t}-\Gamma_{t},v_{t}-\Gamma_{t})\|^{p}_{H^{\eta^{\prime}-1}}\,\mathrm{d}s.$
@@
In the above we used that $H^{\eta}\hookrightarrow B^{\eta^{\prime}+1-2(1+\kappa)/{p}}_{2,p}$ and $B^{\eta^{\prime}+1-2/p}_{2,p}\hookrightarrow H^{\eta^{\prime}}$ (here $B$ stands for Besov spaces), by the above choice of $\kappa$ and $p$.
From the above and Lemma 5.5, we obtain, a.s.,
$\|v_{1}^{\prime}\|_{H^{\eta^{\prime}}}^{p}$ $\textstyle\lesssim\|v\|_{H^{\eta}}^{p}+\sup_{t\in[0,1]}\big{(}\|v_{t}\|_{H^{\eta}}^{2p}+\|\Gamma_{t}\|_{H^{\eta}}^{2p}\big{)}$
$\leq\exp^{[\gamma_{0}]}(R(1+\|v\|_{H^{\eta}}^{\gamma_{1}}+\|(\Gamma_{t})_{t}\|_{L^{\infty}(0,T;H^{\eta})}^{\gamma_{1}})),$
for some $\gamma_{0},\gamma_{1}$ and $R$ depending only on $\sigma$ and $\sigma^{\prime}$. Now, let $\gamma$ be as in (5.9). From the above, we have that, a.s.,
$\log^{[\gamma+\gamma_{0}]}(\|v_{1}\|_{H^{\eta^{\prime}}})\lesssim_{\sigma,\sigma^{\prime}}1+\|(\Gamma_{t})_{t}\|_{L^{\infty}(0,T;H^{\eta})}^{\gamma_{1}}+\log^{[\gamma]}(\|v\|_{H^{\eta}}).$
@@
The implication (5.9$\Rightarrow$(5.10) now follows from taking the expected values, using $\mathbf{E}\|(\Gamma_{t})_{t}\|_{L^{\infty}(0,T;H^{\eta})}^{\gamma_{1}}\leq C(\mathcal{E}_{\sigma})$, and integrating over $\mathbb{H}^{1}_{0}$ with respect to the invariant measure $\mu$. ∎

### v2

@@ 3.2. Lagrangian chaos for stochastic PEs
  • Let $\mu$ be the unique invariant measure for (1.1) on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})$ given by Theorem 3.3. The unique ergodic measure on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$ for the Lagrangian process $(v_{t},\mathbf{x}_{t})$ is given by $\mu\otimes\mathrm{Leb}$.
  • The projective process $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ has a unique ergodic measure on $\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}\times\mathrm{Pr}_{2}$, and it is the unique element of the set
    $\big{\{}\nu\text{ is an invariant measure of $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ satisfying $\nu(\cdot\times\mathrm{Pr}_{2})=\mu\otimes\mathrm{Leb}$}\big{\}}.$
    The same holds with $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ replaced by $(v_{t},\mathbf{x}_{t},\widecheck{\xi}_{t})_{t}$.
Proof.
We first consider the Lagrangian process. The invariance of $\mu\otimes\mathrm{Leb}$ is immediate due to $\nabla\cdot u_{t}=0$. Moreover, due to Proposition 3.11, $\mu\otimes\mathrm{Leb}$ is also ergodicity due to Lemma A.1 (or via Corollary A.2 and Theorem 3.3). The uniqueness part now follows from a well-known variant of Khasminskii’s uniqueness result (see, for example, [25, Proposition 4.1.1]) and Propositions 3.11-3.12.
Second, we consider the projective process $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$, the other being similar. As above, the uniqueness of ergodic measures of $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ follows from Propositions 3.11-3.12 and the above-mentioned variant of the Khasminskii theorem. Let $\mathscr{Q}$ be the set of invariant measures of $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ satisfying $\nu(\cdot\times\mathrm{Pr}_{2})=\mu\otimes\mathrm{Leb}$. This set is not empty due to the Krylov-Bogoliubov procedure, cf., [16, Lemma 3.15] with $\mu$ replaced by $\mu\otimes\mathrm{Leb}$. By the Krein-Milman and Prokhorov theorems, the convex set $\mathscr{Q}$ has extreme points, and $\mathscr{Q}$ coincides with the closed convex hull of its extreme points. By Lemma A.3 and the first part of this corollary, it follows that the extreme points of $\mathscr{Q}$ are ergodic measures. Hence, an ergodic measure for $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ exists, is unique among ergodic measures, and is the unique extreme point of $\mathscr{Q}$. ∎
@@
We begin by proving (3.4) (resp., (3.5)). The latter follows from Propositions 3.8-3.10 and Theorem 3.3 by applying combining the multiplicative ergodic theorem [11, Theorem 3.4.1] (resp., random multiplicative ergodic theorem [50, Theorem III.1.2]) and [16, Proposition 4.17] to the linear cocycle $\mathcal{A}_{\omega,(v,\mathbf{x})}^{t}$ induced by (1.1). Let us note that multiplicative ergodic theorems ensure the existence of the limits (3.4) and (3.5) only for $\mu\otimes\mathrm{Leb}$-a.a. $(v,x)\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$ (here we also used Theorem 3.3). As noticed in [16, Remark 2.2], the strong Feller property as in Proposition 3.11 allow to upgrade the statement to for all $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$. For the reader’s convenience, we include some details. We only comment on the case (3.4) as the other is similar. Suppose that (3.4) holds for a.a. $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$. Let $\mathcal{N}\stackrel{{\scriptstyle{\rm def}}}{{=}}\{(\omega,v,\mathbf{x})\in\Omega\times\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}\,:\,\lim_{t\to\infty}t^{-1}\log|\nabla\phi^{t}_{\omega,v}(\mathbf{x})|=\lambda\}$ and set
$\psi_{\mathcal{N}}(v,\mathbf{x})\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathbf{E}[\mathbf{1}_{\mathcal{N}}(\cdot,v,\mathbf{x})]=\mathbf{E}[\mathbf{1}_{\mathcal{N}}(\theta^{t}\cdot,v,\mathbf{x})]\ \ \text{ for all }t>0,$
@@
where the last equality follows from $\mathbf{P}\cdot(\theta^{t})^{-1}=\mathbf{P}$. From the assumption it follows that $\psi_{\mathcal{N}}(v,\mathbf{x})=1$ for $\mu\otimes\mathrm{Leb}$-a.a. $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$. By the latter and invariance of $\mu\otimes\mathrm{Leb}$ for the Lagrangian process (with $L_{t}$ as in (3.22)),
${1=\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}}\psi_{\mathcal{N}}(v,\mathbf{x})\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}=\int_{\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathbb{T}^{3}}L_{t}\psi_{\mathcal{N}}(v,\mathbf{x})\,\mathrm{d}\mu(v)\,\mathrm{d}\mathbf{x}}$
@@
for all $t>0$ (in the following, we pick $t=1$). Hence, $L_{1}\psi_{\mathcal{N}}(v,\mathbf{x})=1$ for $\mu\otimes\mathrm{Leb}$-a.a. $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$, and by strong Feller property of Proposition 3.11, we have $L_{1}\psi_{\mathcal{N}}(v,\mathbf{x})=1$ for all $(v,\mathbf{x})\in\mathbb{H}^{\sigma}_{0}(\mathbb{T}^{3})\times\mathcal{O}$. Now, the independence of $\theta^{1}\omega$ and $(u_{1}(\omega),\mathbf{x}_{1}(\omega))$ yield
${\int_{\Omega}\mathbf{1}_{\mathcal{N}}(\theta^{1}\omega,v_{1}(\omega),\mathbf{x}_{1}(\omega))\,\mathrm{d}\mathbf{P}(\omega)=L_{1}\psi_{\mathcal{N}}(v,\mathbf{x})=1}.$
@@
Thus, $\mathbf{1}_{\mathcal{N}}(\theta^{1}\cdot,v_{1},\mathbf{x}_{1})$ a.s. Hence, the conclusion follows by noticing that $(\omega,v,\mathbf{x})\in\mathcal{N}$ if and only if $(\theta^{1}\omega,v_{1}(\omega),\mathbf{x}_{1}(\omega))\in\mathcal{N}$ due to the invertibility of $\nabla\phi^{t}_{\cdot,\cdot}(\cdot)$.
@@ 4. Integrability of invariant measures
In this section, we prove Proposition 3.9. As discussed in Subsection 3.3, Proposition 3.9 was used to obtain the existence of the (top) Lyapunov exponent in Theorems 1.1 and 3.5. Here, we prove a more refined result compared to Proposition 3.9 under minimal requirements on $Q$. More precisely, in this subsection, we work under the following assumption.
@@
The following stronger version of Proposition 3.9 is the main result of this section.
@@
(4.1) $\int_{\mathbb{H}_{0}^{1}(\mathbb{T}^{3})}\int_{\mathcal{O}}|\nabla u(\mathbf{x})|^{p}\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)\leq C(p,\mathcal{E})$
@@
(4.2) $\int_{\mathbb{H}_{0}^{1}(\mathbb{T}^{3})}\int_{\mathcal{O}}|\nabla^{2}v(\mathbf{x})|\,\mathrm{d}\mathbf{x}\,\mathrm{d}\mu(v)$
@@
$\lesssim\|\nabla v\|_{L^{2}(\mathbb{T}^{3})}^{2}.$
The last term can be estimated by means of the energy inequality, i.e., Proposition 4.3. From the estimate (4.7), one sees that we have at our disposal more integrability in the vertical direction than in the horizontal ones. The above inequality, along with Proposition 4.3, suggests the validity of (4.5) with $p$ close to $1$.
@@ 4.1. Auxiliary results
In this subsection, we collect some useful auxiliary results. We begin with an elementary lemma concerning the integral operator on the one-dimensional torus $\mathbb{T}=[0,1]$. For $f\in L^{1}(\mathbb{T})$, let
@@
In the proof below, we use the notion of the $H^{\infty}$-calculus and $\mathcal{R}$-sectoriality for various versions of the Laplace operator. We refer to either [48, Chapter 10] or [67, Chapters 3 and 4] for details and notation.
We begin by proving (4.10). Let $A\stackrel{{\scriptstyle{\rm def}}}{{=}}-\Delta_{x,y}$ and $B\stackrel{{\scriptstyle{\rm def}}}{{=}}-\partial^{2}_{z}$ be the realisation of the (minus) Laplace operator on $H^{s_{0},p_{0}}_{x,y}(H^{s_{1},p_{1}}_{z})$ with domain
@@
Moreover, $A$ and $B$ have $H^{\infty}$-calculus of angle 0 by the periodic version of [48, Theorem 10.2.25]. In particular the operator $A+B$ on $\mathsf{D}(A)\cap\mathsf{D}(B)$ is $\mathcal{R}$-sectorial of angle 00 by [67, Corollary 4.5.9]. The maximal $L^{q}$-regularity estimate (4.10) now follows from the Weis’ characterization, see e.g. [67, Theorem 4.4.4] or [72].
Due to (4.10), the semigroup $(e^{t\Delta})_{t\geq 0}$ is analytic on $\mathcal{X}$ by [67, Proposition 3.5.2(i)]. Hence (4.11) follows from Young’s convolution inequality.
It remains to prove the last assertion of Lemma 4.6. The embedding essentially uses mixed-derivative estimates. Indeed, by [67, Corollary 4.5.11] for all $\theta\in(0,1)$,
@@
Before going into the proof of Proposition 4.4, let us state the following result that is a straightforward consequence of the definitions in (2.5) and will be used below without further mention.
@@
The above is known to experts. However, for the reader’s convenience, we include a short proof based on a duality argument. Note that the condition $-d>r-\frac{d}{q}$ is equivalent to $-r-\frac{d}{q^{\prime}}>0$ where $q^{\prime}=\frac{q}{q-1}$. By the vector-valued Sobolev embeddings (see e.g. [61, Proposition 7.4]) we have
@@ 5.1.1. Sobolev estimates for the PEs nonlinearity
Let $\eta,\xi$ be stopping times such that $0\leq\eta\leq\xi\leq\tau_{j}$ a.s. for some $j\geq 1$. By well-known $L^{2}$-estimates for the stochastic heat equation, there exists $C_{0}>0$ independent of $(j,\eta,\xi)$ such that (see e.g., [10, Lemma 4.1] or [60, Chapter 4])
@@
Recall that $\mathcal{E}_{\sigma}=\|Q\|_{\mathscr{L}_{2}(\mathbb{L}^{2}_{0},\mathbb{H}^{\sigma}_{0})}$ and fix $\sigma^{\prime}<1+\sigma$. We claim that, if there exists $\eta\in[2,\sigma]$ and $\gamma\in{\mathbb{N}}_{\geq 1}$ for which
@@
then there exists $\gamma^{\prime}$ depending only on $\gamma,\eta,\sigma$ and $\sigma^{\prime}$ such that
@@
(5.10) $\int_{\mathbb{H}^{1}_{0}(\mathbb{T}^{3})}\log^{[\gamma^{\prime}]}(R_{\gamma}^{\prime}(1+\|v\|_{H^{\eta^{\prime}}}))\,\mathrm{d}\mu(v)\lesssim_{\mathcal{E}_{\sigma}}1\ \ \ \text{ where }\ \ \ \eta^{\prime}\stackrel{{\scriptstyle{\rm def}}}{{=}}\eta+\sigma^{\prime}-\sigma,$
and $R^{\prime}_{\gamma}$ depends only on $\gamma,\eta,\sigma$ and $\sigma^{\prime}$.
To prove (5.9$\Rightarrow$(5.10), we employ weighted maximal $L^{p}$-regularity estimates, see e.g., [67, Section 3.5.2] or [46, Subsection 17.2.e]. Recall that, for
$\eta_{0}\in(\eta^{\prime},\eta+1),$
the operator on $-\Delta:H^{\eta_{0}}\subseteq H^{\eta_{0}-2}\to H^{\eta_{0}-2}$ has weighted maximal $L^{p}$-regularity estimates (this follows from the periodic version of [48, Example 10.1.5]). As we will see below, we do not take $\eta_{0}$ equal to $\eta+1$ due to the $\delta$-loss in Lemma 5.2(3).
@@
Next, let us fix $p\in(2,\infty)$ such that $\eta_{0}-2/p>\eta^{\prime}$, and afterwards $\kappa\in[0,p-1)$ such that $\eta_{0}-2\frac{1+\kappa}{p}<\eta$. Note that the latter choice is always possible by using $\eta^{\prime}<\eta+\frac{1}{2}$ and choosing $\kappa$ close to $p-1$. Hence, by either [46, Corollary 17.2.37] or [67, Theorem 3.4.8], the maximal $L^{p}$-regularity of $-\Delta$ on $H^{\eta_{0}-2}$ with time weight $t^{\kappa}$ applied to $v^{\prime}_{t}=v_{t}-\Gamma_{t}$ (here $\Gamma_{t}=\int_{0}^{t}e^{\Delta(t-s)}Q\,\mathrm{d}W_{s}$), a.s., it holds that
@@
$\|v_{1}^{\prime}\|_{H^{\eta^{\prime}}}^{p}$ $\textstyle\stackrel{{\scriptstyle(i)}}{{\lesssim}}\|v\|_{H^{\eta}}^{p}+\int_{0}^{1}t^{\kappa}\|b(v_{t}-\Gamma_{t},v_{t}-\Gamma_{t})\|^{p}_{H^{\eta_{0}-2}}\,\mathrm{d}s$
$\textstyle\stackrel{{\scriptstyle(ii)}}{{\lesssim}}\|v\|_{H^{\eta}}^{p}+\sup_{t\in[0,1]}\big{(}\|v_{t}\|_{H^{\eta}}^{2p}+\|\Gamma_{t}\|_{H^{\eta}}^{2p}\big{)},$
@@
where in $(i)$ we used $H^{\eta}\hookrightarrow B^{\eta_{0}-2(1+\kappa)/{p}}_{2,p}$ and $B^{\eta_{0}-2/p}_{2,p}\hookrightarrow H^{\eta^{\prime}}$ by the above choice of $\kappa$ and $p$, and in $(ii)$ Lemma 5.2(3) and $\eta_{0}<\eta+1$. As usual, $B$ stands for Besov spaces. From the above and Lemma 5.5, we obtain, a.s.,
$\|v_{1}^{\prime}\|_{H^{\eta^{\prime}}}^{p}\leq\exp^{[\gamma_{0}]}(R(1+\|v\|_{H^{\eta}}^{\gamma_{1}}+\|(\Gamma_{t})_{t}\|_{L^{\infty}(0,1;H^{\eta})}^{\gamma_{1}})),$
for some $\gamma_{0},\gamma_{1}$ and $R$ depending only on $\sigma$ and $\sigma^{\prime}$.
Now, let $\gamma$ be as in (5.9). From the previously displayed formula, we have, a.s.,
$\log^{[\gamma+\gamma_{0}]}(\|v_{1}\|_{H^{\eta^{\prime}}})\lesssim_{\sigma,\sigma^{\prime}}1+\|\Gamma_{1}\|_{H^{\sigma^{\prime}}}+\|(\Gamma_{t})_{t}\|_{L^{\infty}(0,1;H^{\eta})}^{\gamma_{1}}+\log^{[\gamma]}(\|v\|_{H^{\eta}}).$
By stochastic maximal $L^{p}$-regularity for the Laplacian on $H^{\sigma-1}(\mathbb{T}^{3})$ (with no time weight) and $p(\sigma,\sigma^{\prime})<\infty$ such that $\sigma^{\prime}<\sigma+1-\frac{2}{p}$ (see e.g., [7, Theorem 7.16] and [48, Example 10.1.5]), it follows that
$\mathbf{E}\|\Gamma_{1}\|_{H^{\sigma^{\prime}}}^{p}\lesssim_{p,\mathcal{E}_{\sigma}}1.$
@@
Hence, the implication (5.9$\Rightarrow$(5.10) now follows from taking the expected values, using the standard estimate $\mathbf{E}\|(\Gamma_{t})_{t}\|_{L^{\infty}(0,T;H^{\eta})}^{\gamma_{1}}\lesssim_{\mathcal{E}_{\sigma}}1$ (here we used $\eta\leq\sigma$), and integrating over $\mathbb{H}^{1}_{0}$ with respect to the invariant measure $\mu$. ∎